OpenStudy (anonymous):

3 moles of H_2 and 4 moles of I_2 were heated in a sealed tube at a given temperature at which Kc is 50. If the volume of the tube is 1 dm^3 , determine the composition of the equilibrium. H_2 + I_2 ⇌ 2HI

OpenStudy (anonymous):

help plzz!! its been 3 hours !!

OpenStudy (jfraser):

what's the form for the equilibrium expression for this reaction?

OpenStudy (anonymous):

product over reactants

OpenStudy (jfraser):

what about the equation? K = ?? for this reaction?

OpenStudy (anonymous):

i just need help on the table !! rest i can do

OpenStudy (anonymous):

K=50

OpenStudy (jfraser):

what's the equation for K for this reaction?

OpenStudy (anonymous):

\[Kc=\frac{ [HI]^2 }{ [H_2][I_2] }\]

OpenStudy (jfraser):

that's the one

OpenStudy (jfraser):

You know you start with 3 moles of H2 and 4 moles of I2, and you know the equilibrium value has to be 50, so plug into the equation for Kc and solve for the equilibrium concentrations

OpenStudy (anonymous):

so could u help me in the table ?

OpenStudy (anonymous):

H2 + I2 --> 2HI Moles at initial state = 3 4 0

OpenStudy (anonymous):

Moles at equilibrium state ??

OpenStudy (jfraser):

R \(H_2 + I_2 \rightleftharpoons 2HI\) I 3M 4M "0" C "-x" "-x" "+2x" E what are the equilibrium values going to be?

OpenStudy (anonymous):

thats what i need to know

OpenStudy (anonymous):

ok i'll try

OpenStudy (anonymous):

H2 + I2 --> 2HI moles at e = 3-x 4-x 7-x

OpenStudy (anonymous):

is this correct ? coz i m not sure about this

OpenStudy (anonymous):

what happened ??

OpenStudy (anonymous):

hello u there ?

OpenStudy (jfraser):

the equilibrium concentration of HI is 2x, since you start with nothing, and gain twice the amount that the H2 loses. Look at the coefficients of the balanced reaction

OpenStudy (anonymous):

u mean 7-2x

OpenStudy (jfraser):

no, i mean 2x

OpenStudy (jfraser):

where does the 7 come from? from adding the H2 and I2 together? why would you do that?

OpenStudy (anonymous):

oh i was mistaken

OpenStudy (anonymous):

so this is the equation

OpenStudy (anonymous):

\[Kc=\frac{ [2x]^2 }{ [3-x][4-x] }\]

OpenStudy (jfraser):

yes, and the whole value of Kc is 50.

OpenStudy (jfraser):

so solve for X, and you'll get the equilibrium concentrations

OpenStudy (anonymous):

for all 3 ?

OpenStudy (jfraser):

solve for x, then plug each x back into the equilibrium concentration of each species

OpenStudy (anonymous):

oh yes ..i remember !! thanks

OpenStudy (jfraser):

yw

OpenStudy (anonymous):

x^2-35x+600=0

OpenStudy (anonymous):

sorry it is 46x^2-350x+600=0

OpenStudy (jfraser):

then solve for x

OpenStudy (anonymous):

which value should i use for eq. conc. x=5 or x=60/23 ??

OpenStudy (jfraser):

can the equilibrium concentration of H2 be LARGER thanthe amount of H2 you start with?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x=60/23 ? right ?

OpenStudy (jfraser):

so which solution for x is the only one that makes sense?

OpenStudy (anonymous):

x=60/23

OpenStudy (jfraser):

right, now find the equilibrium concentration of H2

OpenStudy (anonymous):

0.391

OpenStudy (jfraser):

and the concentrations of the rest of the reaction

OpenStudy (anonymous):

I2 = 1.391

OpenStudy (anonymous):

and 2HI = 5.217

OpenStudy (jfraser):

look right to me. to check, plug those values into K and make sure you get 50

OpenStudy (anonymous):

YES IT IS CORRECT IT IS COMING 50.04